# Observation of the quantum equivalence principle for matter-waves

### v2

@@ III Results
In Fig. 2 we present the raw data, where we observe about 13 oscillations, which constitute about 80 rad phase accumulation. In Fig. 3 we analyze the phase extracted from Fig. 2, and compare it to a numerical simulation, the analytical prediction of Eq. (4) as well as to the Gedanken experiment [Eq. (2)]. The analytical approximation assumes square pulses of a homogeneous gradient along the $z$-axis giving rise to a purely one-dimensional motion along $z$, which together with the symmetry of the scheme results in perfect path recombination. The numerical simulation takes into account (i) three-dimensional propagation in a curved potential resulting in a non-perfect path recombination, (ii) internal wave packet dynamics (expansion, focusing and rotation) and (iii) atom-atom interactions within each wave packet. The first two factors may change the phase by an order of 1-2 radians each, while the latter may contribute up to about 0.25 radians. Obviously, the prediction of the numerical simulation is accurate to within the uncertainty limits of the experimental parameters. In the SM, we provide detailed information as to how all lines in Figs. 2 and 3 were plotted. The main source of uncertainty in the experimental results arises from the systematic uncertainty in the ratio of the kick and holding currents, as the phase accumulation is most sensitive to this ratio, where deviations from the desired ratio add a phase term that scales linearly with the interferometer time.
@@ IV Discussion
We have realized a unique interferometer, the QGI. While the Ramsey-Bordé interferometer is mostly characterized by the constant spatial separation between the two interferometer paths [53], and the Kasevich-Chu interferometer is mostly characterized by the piece-wise constant velocity separation (difference) between the two paths [54], the QGI enables the control of all degrees of freedom, obtaining spatial, velocity and acceleration differences, almost at will. This allowed us to have one wave packet at rest and the other free-falling.
The cubic phase holds special meaning in the case of our interferometer. In this case it is equal to the phase constituting the quantum equivalence principle or the phase of the transformation between the Newtonian and Einsteinian frames of reference. Another point of view explaining how the QGI enables to probe this fundamental phase lies in the fact that it gives rise to a perfect symmetry between the two frames, as exemplified by the calculation of the actions as seen from both frames. The QGI is perfectly suited for its mission also because it does not require any transformation of electromagnetic waves, so that its physics is only a product of de-Broglie waves.
The good agreement between the data and the analytical as well as numerical prediction, brings us to conclude that we have indeed observed the phase accumulation of a free-falling object, and in doing so, we have confirmed with a high level of confidence the expected $T^{3}$-phase behavior. We have thus observed the phase associated with the equivalence principle in the quantum domain.
As an outlook let us note two important points: First, as our unique interferometer is also able to manipulate clock wave packets (for clocks in a SGI, see  [55]), it opens the door for the test of the more complex formulations suggested for the equivalence principle in the quantum domain. Second, as noted in the following works, such an interferometer also opens the door to searches for new physics [56]. Furthermore, such an interferometer is a milestone towards a SGI with nano-diamonds instead of atoms, for a test of quantum mechanics in new regimes and to probe the quantum-gravity interface, e.g., test the quantization of gravity [49]. In addition, such a large-mass interferometer can explore the Díosi-Penrose conjecture regarding gravitationally-induced collapse, and possible tensions which may arise with the quantum equivalence principle measured in this work (SM).
@@ V Acknowledgments
We thank the BGU nano-fabrication facility for the high-quality chip and the BGU support team, especially Menachem Givon, Zina Binstock, Dmitrii Kapusta and Yaniv Bar-Haim for their support in building and maintaining the experiment. Funding: This work was funded, in part, by the Israel Science Foundation (grants no. 856/18, 1314/19, 3515/20, and 3470/21), and the German-Israeli DIP project (Hybrid devices: FO 703/2-1) supported by the DFG. Additional funding was provided by the Gordon and Betty Moore (doi.org/10.37807/GBMF11936), and Simons (MP-TMPS-00005477) foundations. IF thanks an anonymous USA philanthropist, John Moussouris, Jussi Westergren and the Emmy Network for support and research funding. CM and VV thank the Gordon and Betty Moore Foundation and the Templeton Foundation and CM also thanks the Eutopia Foundation, and Ben Vass for supporting her research. W.P.S. is grateful to the Hagler Institute for Advanced Study at Texas A&\&&M University for a Faculty Fellowship, and to Texas A&\&&M AgriLife Research for the support of this work. We thank Časlav Brukner, Daniel Rohrlich, Daniel M. Greenberger, Ernst M. Rasel, Yair Margalit, J. Ströhle, Gary G. Rozenman, Fabio Di Pumpo and Matthias Zimmermann for helpful discussions.
@@ VI Methods
Let us describe the experimental sequence in some detail (see Fig. 4): we create a 87Rb BEC in the state $|2\rangle\equiv|F=2,m_{F}=2\rangle$ in a magnetic trap produced by current-carrying wires on the atom chip (below the chip which is upside down). We then release the cloud of atoms from the trap and conduct Delta-kick cooling (DKC), which collimates the expansion of the wave packet along the z axis down to an effective temperature of 3.3 nK. As a side effect of this procedure, the atoms are launched upwards (as can be seen on the left side of Fig. 4). The BEC is by now dilute enough so that atom-atom interaction is minimized and the physics becomes single-atom physics. We then apply $100\,\mu$s RF pulse transferring atoms to the state $|F=2,m_{F}=1\rangle$. With the help of a $\pi/2$ pulse, each atom is now in a superposition of a magnetic sensitive state $|1\rangle\equiv|F=2,m_{F}=1\rangle$ and a non-sensitive state $|0\rangle\equiv|F=1,m_{F}=0\rangle$. We now apply a magnetic kick which launches the sensitive state upwards. Quickly thereafter, we apply a $\pi$ pulse, so that the atoms launched upwards are no longer affected by the magnetic gradients and they go into a ballistic trajectory affected only by gravity (up to second order Zeeman). Conversely, the other wave packet becomes sensitive to the magnetic gradient which is now tuned to exactly counteract gravity, so that the wave packet remains stationary in the frame of reference of the laboratory. Finally, we reverse the sequence – apply a $\pi$ pulse and apply the same SG kick to stop the ballistic wave packet, thus bringing the two wave packets to have the same position and momentum for a full overlap. After one more MW $\pi/2$ pulse, another magnetic gradient kick separates the two spin states in space, so that after some time of flight we are able to determine the population in each output port and consequently in each spin state, and thus measure the output of the interferometer (Fig. 2).
@@
(A) The complete experimental scheme of our longitudinal (1D) QGI, for the case of $2T=2.4\,\rm ms$. The x-axis is time, where we set $t=0$ at the moment of trap release. At this time the atoms are in the state $|2\rangle\equiv|F=2,m_{F}=2\rangle$. We apply Delta-kick cooling (DKC) at $t=1.1\,\rm ms$, which collimates the wave packet’s expansion and launches the atoms into a ballistic trajectory upwards. Before the start of the interferometer, we transfer the atoms to the state $|F=2,m_{F}=1\rangle$ by applying an on-resonance RF $\pi$ pulse (grey wavy line). During the interferometer, we control the spin state of the atoms with four MW pulses (orange wavy lines) and the momentum of the atoms with three magnetic gradient pulses (green areas). The MW pulses are on resonance with the transition between the $|1\rangle\equiv|F=2,m_{F}=1\rangle$ and $|0\rangle\equiv|F=1,m_{F}=0\rangle$ states, where the $|0\rangle$ state is magnetically non-sensitive in the Zeeman first order. The first $\pi/2$ pulse puts the atoms in an equal superposition of $|1\rangle+|0\rangle$. At the end of the interferometer, the trajectories of these two states are joined to a single trajectory with two spin states, after which the second $\pi/2$ pulse is applied, creating four overlapping wave packets which interfere, two wave packets per each spin state, imprinting the phase difference into a population difference (for simplicity of graphics, we have not differentiated the four wave packets in the plot). After the second $\pi/2$ pulse, we detect the relative population in each state. In the detection scheme, we first spatially separate the states by applying a fourth magnetic gradient and, finally, image the atoms using on-resonance absorption imaging. The symmetric temporal positioning of the $\pi/2$ and $\pi$ pulses also serves as a dynamic-decoupling scheme, increasing the coherence time, canceling the linear phase accumulation due to possible detuning of the MW and reducing the shot-to-shot phase fluctuations. (B) Zoom-in on the interferometer sequence. The first gradient (kick) pulse applies a maximal acceleration $a_{kick}$ for a total duration of $T_{kick}$, which launches the top arm into a ballistic trajectory. $50\,\rm\mu s$ after the end of the gradient pulse, we apply a $\pi$ pulse (with a duration of $16\,\rm\mu s$) that inverts the spin state of the arms so that the bottom arm is now in the magnetically-sensitive state $|1\rangle$. $5\rm\,\mu s$ after the $\pi$ pulse, we start the holding pulse (duration of $T_{h}$ and rise time $\tau_{h}=12\,\rm\mu s$), which applies an acceleration opposite and equal to gravity, to hold the bottom arm stationary in the lab frame. $50\,\rm\mu s$ after the holding pulse, we again apply a $\pi$ pulse and $5\rm\,\mu s$ later we recombine the trajectories of the two wave packets by applying a magnetic gradient pulse with a maximal acceleration $a_{kick}$ for a total duration of $T_{kick}$, as in the first pulse.
@@
The trajectories of the two arms of the interferometer, calculated by the numerical simulation for holding durations $T_{h}=232,742,1252,1762,2272\,\rm\mu$s. Dashed blue lines represent the ballistic trajectory, and solid red lines the static wave packet. At the end of the interferometer, a small spatial splitting can be seen, on the order of $0.1\,\rm\mu m$. For $T_{h}=2272\,\rm\mu$s, the spatial splitting between the two arms halfway through the interferometer reaches $7.5\,\rm\mu$m. The trajectories are plotted until the time in which the final (second) $\pi/2$ is applied.
@@ VII.7.2 Derivation of the phase difference using the action
$K_{j}(t)=\frac{1}{2}mv_{j}^{2}(t)\;,\;U_{j}(t)=-m(g+a_{j})z_{j}(t)\,.$ (13)
To maintain the symmetry of the trajectories we set the initial conditions of the COM of the wave packet to be
@@
We define the total interferometer time as $2T+2T_{kick}$, and the holding duration of the bottom arm can be defined as $T_{h}=2T-2T_{d}$. The experiment can be described by a set of durations $(T_{kick},T_{d},T_{h},T_{d},T_{kick})$, where the acceleration of the ballistic and reference paths in each of these durations is given by $a_{ballistic}=-g+(a_{kick},0,0,0,a_{kick})$ and $a_{reference}=-g+(0,0,a_{hold},0,0)$. The condition for a closed interferometer is $a_{kick}=\frac{g}{2T_{kick}}T_{h}$ and setting $a_{hold}=g$ confirms that the reference path is stationary in the lab frame. Performing the integrals, we get
@@
$\lim_{T_{kick}\to 0}\Delta\phi=\frac{mg^{2}}{3\hbar}\left(T^{3}-T_{d}^{3}\right).$ (18)
@@
The result of the analytical model is given here in Eq. (17) (and in Eq. (7) of the main text) is plotted in Fig. 3 of the main text using the experimental values of $T_{kick}=80\rm\,\mu s$ and $T_{d}=71+6\rm\,\mu s$ where we include half the rise time of the holding pulse $\tau_{\rm hold}/2=6\rm\,\mu s$ in the delay. In the analytical model prediction, we also include the effect of the second-order Zeeman energy of the $|0\rangle$ state. This energy results in a small force on the $|0\rangle$ state during the gradient pulses, approximately $1\%$ of the force on the $|1\rangle$ state, and in the opposite direction. To account for this second-order effect, we calculate the acceleration of the $|0\rangle$ state due to second order Zeeman effect during the holding stage $a_{\rm SOZ}=0.10\rm\,m/s^{2}$ (see Sec. K for calculation of $a_{\rm SOZ}$). We set $g_{eff}=g+a_{\rm SOZ}=9.91\,\rm m/s^{2}$ in the analytical prediction. Accounting to $a_{\rm SOZ}$ by setting $g_{eff}=g+a_{\rm SOZ}$ is accurate up to the leading order of $T$, and comparing the analytical model that includes $a_{\rm SOZ}$ explicitly to setting $g_{eff}=g+a_{\rm SOZ}$ shows that the two agree up to $0.03\,\rm rad$.
The Gedanken experiment is the case of the simple thought experiment, in which the kick pulse is a delta pulse so that $T_{kick}=0$, and the technical delay between the pulses vanishes setting $T_{d}=0$. The prediction is given by setting $T_{kick}=0$ and $T_{d}=0$ in Eq. 15 (or 6) which results in Eq. 19.

### v3

@@ III Results
In Fig. 2 we present the raw data, where we observe about 13 oscillations, which constitute about 80 rad phase accumulation. In Fig. 3 we analyze the phase extracted from Fig. 2, and compare it to a numerical simulation, the analytical prediction of Eq. (4) as well as to the Gedanken experiment [Eq. (2)]. The analytical approximation assumes square pulses of a homogeneous gradient along the $z$-axis giving rise to a purely one-dimensional motion along $z$, which together with the symmetry of the scheme results in perfect path recombination. The numerical simulation takes into account (i) three-dimensional propagation in a curved potential resulting in a non-perfect path recombination, (ii) internal WP dynamics (expansion, focusing and rotation) and (iii) atom-atom interactions within each WP. The first two factors may change the phase by an order of 1-2 radians each, while the latter may contribute up to about 0.25 radians. Obviously, the prediction of the numerical simulation is accurate to within the uncertainty limits of the experimental parameters. In the SM, we provide detailed information as to how all lines in Figs. 2 and 3 were plotted. The main source of uncertainty in the experimental results arises from the systematic uncertainty in the ratio of the kick and holding currents, as the phase accumulation is most sensitive to this ratio, where deviations from the desired ratio add a phase term that scales linearly with the interferometer time.
@@ IV Discussion
We have realized a unique interferometer, the QGI. While the Ramsey-Bordé interferometer is mostly characterized by the constant spatial separation between the two interferometer paths [55], and the Kasevich-Chu interferometer is mostly characterized by the piece-wise constant velocity separation (difference) between the two paths [56], the QGI enables the control of all degrees of freedom, obtaining spatial, velocity and acceleration differences, almost at will. This allowed us to have one WP at rest and the other free-falling.
The cubic phase holds special meaning in the case of our interferometer. In this case it is equal to the phase constituting the QEP or the phase of the transformation between the Newtonian and Einsteinian frames of reference. Another point of view explaining how the QGI enables to probe this fundamental phase lies in the fact that it gives rise to a perfect symmetry between the two frames, as exemplified by the calculation of the actions as seen from both frames. The QGI is perfectly suited for its mission also because it does not require any transformation of electromagnetic waves, so that its physics is only a product of de-Broglie waves.
The good agreement between the data and the analytical as well as numerical prediction, brings us to conclude that we have indeed observed the phase accumulation of a free-falling object, and in doing so, we have confirmed with a high level of confidence the expected $T^{3}$-phase behavior. We have thus observed the phase associated with the EP in the quantum domain.
As an outlook let us note two important points: First, as our unique interferometer is also able to manipulate clock WPs (for clocks in a SGI, see  [57]), it opens the door for the test of the more complex formulations suggested for the EP in the quantum domain  [13, 14, 15, 16, 17, 18, 19, 20, 58]. Such an interferometer also opens the door to searches for new physics with clocks [59]. Second, such an interferometer is a milestone towards a SGI with nano-diamonds instead of atoms, for a test of quantum mechanics in new regimes and to probe the quantum-gravity interface [60], e.g., test the quantization of gravity [51]. In addition, such a large-mass (active-mass) interferometer can explore the Díosi-Penrose conjecture regarding gravitationally-induced collapse, and possible tensions which may arise with the QEP measured in this work [61, 62] (see SM).
@@ V Acknowledgments
We thank the BGU nano-fabrication facility for the high-quality chip and the BGU support team, especially Menachem Givon, Zina Binstock, Dmitrii Kapusta and Yaniv Bar-Haim for their support in building and maintaining the experiment. Funding: This work was funded, in part, by the Israel Science Foundation (grants no. 856/18, 1314/19, 3515/20, and 3470/21), and the German-Israeli DIP project (Hybrid devices: FO 703/2-1) supported by the DFG. Additional funding was provided by the Gordon and Betty Moore (doi.org/10.37807/GBMF11936), and Simons (MP-TMPS-00005477) foundations. I.F. thanks an anonymous USA philanthropist, John Moussouris, Jussi Westergren and the Emmy Network for support and research funding. C.M. and V.V. thank the Gordon and Betty Moore Foundation and the Templeton Foundation and C.M. also thanks the Eutopia Foundation, and Ben Vass for supporting her research. W.P.S. is grateful to the Hagler Institute for Advanced Study at Texas A&\&&M University for a Faculty Fellowship, and to Texas A&\&&M AgriLife Research for the support of this work. We thank Časlav Brukner, Daniel Rohrlich, Daniel M. Greenberger, Ernst M. Rasel, Yair Margalit, Jannik Ströhle, Gary G. Rozenman, Fabio Di Pumpo and Matthias Zimmermann for helpful discussions.
@@ VI Methods
Let us describe the experimental sequence in some detail (see Fig. 4): we create a 87Rb BEC in the state $|2\rangle\equiv|F=2,m_{F}=2\rangle$ in a magnetic trap produced by current-carrying wires on the atom chip (below the chip which is upside down). We then release the cloud of atoms from the trap and conduct Delta-kick cooling (DKC), which collimates the expansion of the WP along the z axis down to an effective temperature of 3.3 nK. As a side effect of this procedure, the atoms are launched upwards (as can be seen on the left side of Fig. 4). The BEC is by now dilute enough so that atom-atom interaction is minimized and the physics becomes single-atom physics. We then apply $100\,\mu$s radio-frequency (RF) $\pi$ pulse transferring atoms to the state $|F=2,m_{F}=1\rangle$. With the help of a microwave (MW) $\pi/2$ pulse, each atom is now in a superposition of a magnetic sensitive state $|1\rangle\equiv|F=2,m_{F}=1\rangle$ and a non-sensitive state $|0\rangle\equiv|F=1,m_{F}=0\rangle$. We now apply a magnetic SG kick which launches the sensitive state upwards. Quickly thereafter, we apply a MW $\pi$ pulse, so that the atoms launched upwards are no longer affected by the magnetic gradients and they go into a ballistic trajectory affected only by gravity (up to second order Zeeman). Conversely, the other WP becomes sensitive to the magnetic gradient which is now tuned to exactly counteract gravity, so that the WP remains stationary in the frame of reference of the laboratory. Finally, we reverse the sequence – apply a MW $\pi$ pulse and apply the same SG kick to stop the ballistic WP, thus bringing the two WPs to have the same position and momentum for a full overlap. After one more MW $\pi/2$ pulse, another magnetic gradient kick separates the two spin states in space, so that after some time of flight we are able to determine the population in each output port and consequently in each spin state, and thus measure the output of the interferometer (Fig. 2).
@@
(A) The complete experimental scheme of our longitudinal (1D) QGI, for the case of $2T=2.4\,\rm ms$. The x-axis is time, where we set $t=0$ at the moment of trap release. At this time the atoms are in the state $|2\rangle\equiv|F=2,m_{F}=2\rangle$. We apply Delta-kick cooling (DKC) at $t=1.1\,\rm ms$, which collimates the WP’s expansion and launches the atoms into a ballistic trajectory upwards. Before the start of the interferometer, we transfer the atoms to the state $|F=2,m_{F}=1\rangle$ by applying an on-resonance RF $\pi$ pulse (grey wavy line). During the interferometer, we control the spin state of the atoms with four MW pulses (orange wavy lines) and the momentum of the atoms with three magnetic gradient pulses (green areas). The MW pulses are on resonance with the transition between the $|1\rangle\equiv|F=2,m_{F}=1\rangle$ and $|0\rangle\equiv|F=1,m_{F}=0\rangle$ states, where the $|0\rangle$ state is magnetically non-sensitive in the Zeeman first order. The first $\pi/2$ pulse puts the atoms in an equal superposition of $|1\rangle+|0\rangle$. At the end of the interferometer, the trajectories of these two states are joined to a single trajectory with two spin states, after which the second $\pi/2$ pulse is applied, creating four overlapping WPs which interfere, two WPs per each spin state, imprinting the phase difference into a population difference (for simplicity of graphics, we have not differentiated the four WPs in the plot). After the second $\pi/2$ pulse, we detect the relative population in each state. In the detection scheme, we first spatially separate the states by applying a fourth magnetic gradient and, finally, image the atoms using on-resonance absorption imaging. The symmetric temporal positioning of the $\pi/2$ and $\pi$ pulses also serves as a dynamic-decoupling scheme, increasing the coherence time, canceling the linear phase accumulation due to possible detuning of the MW and reducing the shot-to-shot phase fluctuations. (B) Zoom-in on the interferometer sequence. The first gradient (kick) pulse applies a maximal acceleration $a_{kick}$ for a total duration of $T_{kick}$, which launches the top arm into a ballistic trajectory. $50\,\rm\mu s$ after the end of the gradient pulse, we apply a $\pi$ pulse (with a duration of $16\,\rm\mu s$) that inverts the spin state of the arms so that the bottom arm is now in the magnetically-sensitive state $|1\rangle$. $5\rm\,\mu s$ after the $\pi$ pulse, we start the holding pulse (duration of $T_{h}$ and rise time $\tau_{h}=12\,\rm\mu s$), which applies an acceleration opposite and equal to gravity, to hold the bottom arm stationary in the lab frame. $50\,\rm\mu s$ after the holding pulse, we again apply a $\pi$ pulse and $5\rm\,\mu s$ later we recombine the trajectories of the two WPs by applying a magnetic gradient pulse with a maximal acceleration $a_{kick}$ for a total duration of $T_{kick}$, as in the first pulse.
@@
The trajectories of the two arms of the interferometer, calculated by the numerical simulation for holding durations $T_{h}=232,742,1252,1762,2272\,\rm\mu$s. Dashed blue lines represent the ballistic trajectory, and solid red lines the static WP. At the end of the interferometer, a small spatial splitting can be seen, on the order of $0.1\,\rm\mu m$. For $T_{h}=2272\,\rm\mu$s, the spatial splitting between the two arms halfway through the interferometer reaches $7.5\,\rm\mu$m. The trajectories are plotted until the time in which the final (second) $\pi/2$ is applied.
@@ VII.7.2 Derivation of the phase difference using the action
$K_{j}(t)=\frac{1}{2}mv_{j}^{2}(t)\;,\;U_{j}(t)=-ma_{j}(t)[z_{j}(t)-z_{0}]\,,$ (13)
where $z_{0}$ is defined by the zero of magnetic potential, $|B(0,0,z_{0})|=0$, which is the position of the magnetic field quadrupole in the experiment (not to be confused with the initial position of the atoms $z_{i}(0)$). In the case of a closed interferometer, the phase does not depend on $z_{0}$, while for an open interferometer, the phase will depend on $z_{0}$. To maintain the symmetry of the trajectories we set the initial conditions of the COM of the wave packet to be
@@
We define the total interferometer time as $2T+2T_{kick}$, and the holding duration of the bottom arm can be defined as $T_{h}=2T-2T_{d}$. The experiment can be described by a set of durations $(T_{kick},T_{d},T_{h},T_{d},T_{kick})$, where the acceleration of the ballistic and reference paths in each of these durations is given by $a_{ballistic}=-g+(a_{kick},0,0,0,a_{kick})$ and $a_{reference}=-g+(0,0,a_{hold},0,0)$.
The condition for a closed interferometer is an equal area of the accelerations of the two arms,
[…]
which, in this case, gives
$a_{kick}\cdot 2T_{kick}=a_{hold}\cdot T_{h}\,.$ (16)
Setting $a_{hold}=g$ confirms that the reference path is stationary in the lab frame. Performing the integrals, we get
@@
$\lim_{T_{kick}\to 0}\Delta\phi=\frac{mg^{2}}{3\hbar}\left(T^{3}-T_{d}^{3}\right),$ (20)
@@
Eq.21 represents the Gedanken experiment in which the beam splitter is a delta pulse such that $T_{kick},T_{d}\to 0$.
The result of the analytical model given in Eq. (19) (and in Eq. (4) of the main text) is plotted in Fig. 3 of the main text using the experimental values of $T_{kick}=80\rm\,\mu s$ and $T_{d}=71+6\rm\,\mu s$ where we include half the rise time of the holding pulse $\tau_{\rm hold}/2=6\rm\,\mu s$ in the delay. We also include the contribution to the phase from the force on the state $|0\rangle$ due to the second-order Zeeman (SOZ) effect. This force is opposite to the force on the $|1\rangle$ state during the gradient pulses, and its value is about 1% of the latter. During the holding pulse, the acceleration of the state $|0\rangle$ due to SOZ is calculated to be ${a}_{SOZ}=0.10\,\rm m/s^{2}$ (see Sec. K) for details). We include this effect in the analytical prediction for the phase (Eq. 17) by replacing the gravitational acceleration $g$ by $g_{\rm eff}=g+a_{\rm SOZ}=9.91\,\rm m/s^{2}$. This estimation agrees with a model in which $a_{SOZ}$ is explicitly included in the calculation to within $0.03\,\rm rad$.